1.An object moves along a straight with an acceleration of 2m/s2.If its initial speed is 10m/s,what will be its speed 5s later?
2.An object dropped from a cliff falls with a constant acceleration of 10m/s2. Find its speed 2 s after it was dropped?

Answer by Alan3354(30993)

1.An object moves along a straight with an acceleration of 2m/s2.If its initial speed is 10m/s,what will be its speed 5s later?
v = 10 + 2*5 = 20 m/sec
-------------------
2.An object dropped from a cliff falls with a constant acceleration of 10m/s2. Find its speed 2 s after it was dropped?
= 10*2 = 20 m/sec